OpenStudy (anonymous):

If sin angle A=3/5 and cos angle A= 4/5,what is tan angle a A 4/3 B 3/4 C 7/5 D 1/5

OpenStudy (lgbasallote):

sin A = opposite/hypotenuse = 3/5 cos A = adjacent/hypotenuse = 4/5 tan A = opposite/adjacent so what do you think is tan A

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

Thanks

OpenStudy (lgbasallote):

welcome!